# Distribution Matching Distillation Meets Reinforcement Learning

### v3

@@ Appendix A More Related Works
Appendix B Additional Visual Comparison Results
In this section, we provide additional visual comparisons for our 4-step distilled models against teacher models and other methods. The results are shown in Figure 10, Figure 11:
Figure 10: Additional visual comparison in SDXL-Base. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
Figure 11: Additional visual comparison in SD3-Medium. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
Figure 12: Additional visual comparison in SD3.5-Large. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.

### v4

@@ 5 Related Work
Reward models for diffusion post-train. Unlike Large Language Models (LLMs), where reinforcement learning (RL) can often leverage verifiable, rule-based rewards [deepseekmath, deepseekr1, glm] (e.g., code execution success or mathematical correctness), diffusion-based image generation is inherently subjective. The evaluation of generated images relies heavily on multifaceted criteria such as aesthetic quality, instruction following, and so on, which are difficult to quantify through rigid, rule-based functions. Consequently, the development of robust reward models (RMs) has become a critical prerequisite for effective RL-based alignment in diffusion models [refl, rewarddance, clipscore, hps, visionreward, q-align]. Early methods, such as CLIP-score [clipscore], primarily focus on measuring the semantic alignment between the text prompt and the generated image. To address the need for visual fidelity and artistic appeal, metrics like Aesthetic Score [laion5b] and HPS (Human Preference Score) [hps, hpsv2, hpsv3] have been introduced to capture the nuance of human aesthetic judgment. More recently, the reward models become unified [unified-reward, visionreward, unified-p-reward] and aggregate multi-dimensional preferences—ranging from low-level texture details to high-level semantic instruction following—into a unified scoring function. While our primary contribution lies in proposing a novel algorithm that integrates few-step distillation with reinforcement learning, our empirical results highlight the pivotal role of the reward model in our pipeline.
6 Conclusion
We introduces DMDR, a unified framework that transforms Distribution Matching Distillation (DMD) and Reinforcement Learning (RL) from independent, sequential stages into a synergistic training pipeline. We demonstrate that joint optimization is mutually beneficial: RL steers distillation toward high-reward regions to break the ”teacher ceiling,” while DMD regularizes RL to effectively mitigate reward hacking. By employing a two-stage strategy—incorporating RT-DM and dynamic training followed by joint optimization—DMDR achieves state-of-the-art few-step performance across diverse architectures. Ultimately, our approach enables few-step models to not only match but consistently surpass the visual quality and prompt adherence of their original multi-step teacher models.